Astronomy Report – March 17th

Name: Sophie Wuyckens Crew: 190
date: 3/17/2018

Sky Conditions: Clear but sometines some clouds

Wind Conditions: High wind all the day long

Observation Start Time: No observation

Observation End Time: –

Summary: I didn’t dare to use the the observatory today because the wind was very violent.

GreenHab Report – March 17th

GreenHab report for Sol6 :

Crewmember’s name: Frédéric Peyrusson (Biologist)

Date 17/03/2018

Environmental control:

Ambient with door open.

Shade cloth on.

Fan on, without cooling.

Average temperatures: 23°C

Low: 8,4°C

High: 32,8°C

Hours of supplemental light: 0 hours

Changes to crops: No change

Daily water usage for crops: 5 gal

Time(s) of watering for crops: 3

Moringa research observations: ?

Changes to research plants: N/A

Everything went fine, seeds of radish and coriander are now emerging. Mint plants supplemented with hydrogels seem to better tolerate transplantation in various soils than control plants.
All other plants have been watered.

Support/supplies needed: N/A

Journalist Report – March 17th

Apart from each crew member’s experiment, there are many things to do at the MDRS. Cooking, cleaning and the usual chores are of course part of the job, like in any shared living space. But the MDRS being quite a special place to live, as it is, after all, located on Mars, there are additional and essential tasks that need to be completed every day. Taking care of the GreenHab and its plants is an essential one – indeed, they are our life support, producing both our food and our oxygen, and they therefore require exceptional care. However, the MDRS is a complex system, with many variables. If one of them fails, the whole station’s well-being is compromised. The role of the crew engineer, Bastien, is therefore also essential : every day, he checks all the systems, machines, reservoirs, electrical devices, batteries and engines, inside and outside the station, that allow its optimal functioning. His is truly a full-time job. And despite this, Bastien has been involved in quite a serious and time-consuming project of aerial 3D cartography. With his drone, and a computer program designed for this purpose, he has been criss-crossing rectangular areas of terrain, taking dozens of pictures from all angles, which are then automatically recombined in a tedious hours-long process, giving a 3D map of the area, which can then be assembled with surrounding maps to create maps of whole regions.

Michael, our XO and an operations research specialist, has also been working quite hard : his main project concerns something called robust planning under uncertainty, which means that he has been working on a dynamic, adaptable, and predictive schedule for the whole crew. This planning program, coded on his computer, has been designed by Michael from scratch and is still getting improved day by day. Indeed, the point of such a program is to adapt the schedule to unpredicted (and unpredictable) events that slow down or accelerate a particular crew member’s experiment. Since all experiments are related in some way (by equipment, place and/or time), a change in one experiment will often impact many other experiments – this makes for quite a complex system, which needs to be updated daily. Michael has therefore been taking daily notes of our progress, and integrating them into his program to generate the next days’ improved schedule.

As you may think, Michael is quite the computer enthusiast. You can imagine his delight when he realised that the small remote-controlled linux-powered rover called Phoenix, made by members of the NorCal Mars Society, didn’t work and needed troubleshooting. Yesterday evening, he got down to understanding where the problem was, and started working in a trance-like state, measuring voltages, connecting cables and reading and writing code that seems like gibberish to the untrained eye. In the end, he understood that the problem was two-fold : on the one hand, the rover’s internal clock was unsynchronized with the clock of the computer which controlled it, a simple but hard to detect error. On the other hand, there were some battery issues which needed to be resolved, and after 24 hours, the rover was finally set free in front of the base this afternoon, where it performed some test loops.

We are now fully equipped, with Bastien’s drone and NorCal’s rover, to inspect the station from the outside without having to go through the tedious process of getting fully suited up and going through a decompression period in the exit vault. Another step forward for our Martian colony!

EVA #7 Report – March 17th

EVA #7 , Sol6
Crew members: Mario Sundic (EVA leader), Ariane Sablon, Maximilien Richald, Michael Saint-Guillain

Location:
Near the Hab: 518200E-4250750N
Near Zubrin’s head: 518500E-4248500N

Objectives:
Mario: Pictures for Journalist’s Report and photo a a mountain for Sophie’s muography (potential experimental site).
Maximilien: Soil sampling
Ariane: Checking of Martin Roumain’s drug samples previously disposed around the MDRS
Michael: support

Operations:
09:28 entering the sas
09:31 end of depressurization
09:34 departure rovers
09:44 arrival at Zubrin’s head
10:06 arrival at the base of the targeted mountain
10:15 departure form mountain site
10:31 soil sampling
10:37 departure with rovers
10:47 arrival MDRS
10:50 pictures MDRS
10:53 check of Martin’s samples
10:57 entering the sas and pressurization

Narrative:
I am Michael Saint-Guillain, and I am writing this report on behalf of the EVA leader Mario Sundic. We entered the sas in time to start the EVA at 9:30. This was our first EVA in the South of MDRS. It is quite surprising to see how the landscape can be slightly different down there compared to the North sites. I personally totally LOVED the landscapes near Zubrin’s head. Totally Martian, as far as I know. After driving the rovers towards the targeted site, we walked for a rough hour, between the hills, mountains and cliffs… Amazing. I would have spend the entire day there. Unfortunately, the schedule (as well as the objectives) had to be respected, and we proceeded to the various soil samplings and photo shootings (required by Sophie for her project of muon detector). Back at the MDRS, we finally checked Martin’s drug sampling (which he disposed outside the MDRS a couple of days ago) as required and entered the sas, our head filled with Martian landscapes.

Vehicles
Curiosity & Spirit rovers

Sol 6 Summary – March 17th

Crew 190 Sol Summary Report
17March2018
Sol 6

Summary Title:
The Phoenix great escape!

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 6, 08:00

Sol Activity Summary:
8:15 Medic inspection: memory and reflex tests
8:40 Breakfast
9:30 EVA: Mario (photos), Maximilien (soil sampling), Ariane (support), Michael (support)
9:30 Others: Scientific work
11:00 EVA ended; Crew meeting before lunch
12:30 Lunch
14:30 Scientific work: muography (Sophie), cartography of MDRS (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), Phoenix operations & troubleshooting (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan:
Sol 7:

Notes: an EVA is planned in the afternoon, involving Maximilien, Fred, Sophie, Bastien and Mario. Remaining of the day will be dedicated to scientific work.

Sol 7 to Sol 12:

Notes: previsional planning for the second week of the rotation. Very likely to be modified.

Anomalies in work:
We finally get the Phoenix fully working (except the GPS). Unfortunately, Phoenix got trapped in a tiny hole in the soil; we organised (Michael & Bastien) a rescue mission. Phoenix is now back home, a bit scared though.

Weather:
Sunny in the morning. Cloudy in the afternoon. A bit cold.

Crew Physical Status:
A bit tired, hard to wake up.

EVA:
See EVA report.

Reports to be file:
Commander report
Journalist report
EVA#8 request for Sol 7
Engineer report
Pictures

Support Requested:
None

Commander Report – March 17th

Commander Maximilien Richald

We started our day by a breakfast, followed by the daily health test given by Martin the HSO of the crew. Four of us left for an EVA of two hours. We explored the mountain close from Zubrin’s head where we took pictures and harvested one sample of white soil for me.

After eating a potatoes lunch, some of us rested a bit while the rest of the team was working on its experiments. Martin has problem with his calibration curve for his analysis. The turbidity of his solutions was too high to allow a measurement by UV-Vis spectrophotometry. Fortunately, we worked together to find a solution and get a calibration curve of quality.

Michael solved the IT problem with the rover and he was able to make it riding. Unfortunately, the rover, which was riding around the station, got stuck and we organized a rescue mission to bring it back to the Hab.

In the end of the afternoon, Mario cooked cookies for us and everyone took time to relax. Tonight, we will continue our evening with a diner cooked by Ariane and a “life story” night.

At this time of the mission, nobody is complaining about confined environment or difficulties to live on Mars! Let’s continue like this!

Regarding the station, no significant problem was detected.

Operations Report – March 17th

Crew 190 Operations Report 17th March 2018

SOL: 6

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 9.36 am and turned on at 6.10 pm.

Solar – SOC 99% (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~ 250 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 303.3

Toilet tank emptied: NO

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repaired

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: 25.1

Beginning charge: 89%

Ending charge: 100%

Currently charging: YES

Opportunity rover used: NO

Hours: 24.1

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Curiosity rover used: YES

Hours: 19.3

Beginning charge: 86%

Ending charge: 100%

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 250 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Ariane, Sophie & Bastien working on their experiments

Summary of GreenHab operations: –

Summary of ScienceDome operations: Martin, Maximilien & Fred working on their experiments.

Summary of RAMM operations: the muon detector of Sophie is remotely controlled from the Hab

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –